# Internet of Agents: Fundamentals, Applications, and Challenges

### v1

@@ III-F Trustworthy Regulation in IoA
1) DID and Verifiable Credentials (VCs): DID frameworks (e.g., W3C DID [36] and Sovrin [126]) assign agents self-sovereign identifiers, which are independent of centralized authorities and anchored in tamper-resistant ledgers [127]. With VCs, agents can issue, present, and revoke cryptographically signed credentials such as trust ratings, compliance certificates, or capability attestations in a privacy-friendly manner (i.e., without disclosing underlying data) [127]. Modern DID implementations should support on-demand DID creation, context-specific key rotation, and selective disclosure proofs, to enable privacy-preserving attribute validation across domains. However, in time-varying IoA environments, maintaining up-to-date DID documents and revocation statuses at large-scale IoA needs near-instant ledger updates and efficient distributed resolution services, which requires further investigations.
@@
2) Blockchain and Distributed Ledger Technologies. Blockchain (e.g., Hyperledger Fabric and Ethereum) offers an immutable and decentralized infrastructure for managing digital identities, recording agent actions, and enabling transparent auditing [125, 128]. Smart contracts can enforce predefined policies for agent authorization, behavior verification, accountability, and decentralized governance in a trustless environment [129]. Despite its strengths, blockchain suffers from latency, limited throughput, and high energy consumption issues that conflict with real-time coordination in large-scale IoA. Moreover, most blockchains are not designed for privacy-preserving operations [130], which may expose sensitive agent behavior or user data. Emerging solutions such as sharding, off-chain channels, and zero-knowledge proofs mitigate these issues but also add complexity to protocol design and interoperability [131].
@@
3) Regulations and Legal Design. As agents assume decision-making roles in critical domains, comprehensive legal frameworks should define agent owner, liability regimes, and oversight procedures. Regulatory sandboxes [132] and certification processes can vet agent behaviors in controlled environments, while dynamic rule frameworks (combining statutes, industry standards, and ethical guidelines) set operational boundaries. Algorithmic audits and mandated explainability reports further ensure traceability and compliance. Global IoA deployments face jurisdictional fragmentation and conflicting regulations, necessitating harmonized cross-border enforcement protocols and data-sovereignty agreements. Moreover, apportioning liability in multi-agent collaborations, where agents autonomously negotiate and delegate tasks, requires novel legal constructs to assign fault among developers, operators, and the agents themselves, without stifling innovation.
@@ IV Future Research Directions
IV-A Secure & Adaptive Agent Communication Protocols
@@
The design of inter‐agent protocols navigates a versatility-efficiency-portability trilemma [11]: highly versatile protocols handle diverse tasks but incur heavy overhead, while efficient and portable APIs lack flexibility. Agents offer a compromise, using natural language and dynamic tool calls, but at the expense of latency, energy cost, and hallucination risks. Furthermore, as agent teams form and reconfigure dynamically, communication protocols should adapt in real time to changing roles, contexts, and capabilities. Moreover, existing MCP-based ecosystems remain susceptible to tool‐poisoning attacks whereby malicious agents inject compromised plugins or APIs; decentralized A2A networks can suffer Sybil attacks from adversary-controlled identities; and Agora’s reliance on natural-language routines introduces hallucination vulnerabilities when LLM-generated schemas diverge from intended semantics. Future IoA communication should embed rigorous security guarantees to build secure agent ecosystems.
IV-B Decentralized and Self-Governing Agent Ecosystems
@@
IV-C Agent-Based Economic Systems
@@
IV-D Privacy-Preserving Agent Interactions
@@
IV-E Cyber-Physical Secure IoA

### v2

@@ III-F Trustworthy Regulation in IoA
1) Accountability of AI Agents. Accountability in agentic systems requires robust mechanisms for tracing, attributing, and enforcing responsibility across the agent lifecycle. For instance, Chan et al. [137] propose a multi-layered visibility framework, including agent identifiers, real-time monitoring, and activity logging, to enable traceability and post-incident forensics. These measures facilitate the identification of responsible parties, including developers, deployers, or users, when agents cause harm or violate policies. Tamang and Bora [138] introduce the Enforcement Agent (EA) framework, which embeds supervisory agents within multi-agent environments to monitor peer behavior, detect misalignment, and intervene in real time. The results demonstrate that such embedded oversight significantly enhances system safety and distributed accountability, especially in adversarial or dynamic settings. As reported by [139], the “many hands problem” in agent value chains emphasizes the need for clear allocation of accountability among model providers, system providers, and deployers.
@@
  • DID and Verifiable Credentials (VCs). DID frameworks (e.g., W3C DID [36] and Sovrin [140]) assign agents self-sovereign identifiers, which are independent of centralized authorities and anchored in tamper-resistant ledgers [141]. With VCs, agents can issue, present, and revoke cryptographically signed credentials such as trust ratings, compliance certificates, or capability attestations in a privacy-friendly manner (i.e., without disclosing underlying data) [141]. Modern DID implementations should support on-demand DID creation, context-specific key rotation, and selective disclosure proofs, to enable privacy-preserving attribute validation across domains. However, in time-varying IoA environments, maintaining up-to-date DID documents and revocation statuses at large-scale IoA needs near-instant ledger updates and efficient distributed resolution services, which requires further investigations.
@@
  • Blockchain and Distributed Ledger Technologies. Blockchain (e.g., Hyperledger Fabric and Ethereum) offers an immutable and decentralized infrastructure for managing digital identities, recording agent actions, and enabling transparent auditing [136, 142]. Smart contracts can enforce predefined policies for agent authorization, behavior verification, accountability, and decentralized governance in a trustless environment [143]. Despite its strengths, blockchain suffers from latency, limited throughput, and high energy consumption issues that conflict with real-time coordination in large-scale IoA. Moreover, most blockchains are not designed for privacy-preserving operations [144], which may expose sensitive agent behavior or user data. Emerging solutions such as sharding, off-chain channels, and zero-knowledge proofs mitigate these issues but also add complexity to protocol design and interoperability [145].
2) Regulatory Compliance. As agents assume decision-making roles in critical domains, comprehensive legal frameworks should define agent owner, liability regimes, and oversight procedures. Regulatory sandboxes [146] and certification processes can vet agent behaviors in controlled environments, while dynamic rule frameworks (combining statutes, industry standards, and ethical guidelines) set operational boundaries. Algorithmic audits and mandated explainability reports further ensure traceability and compliance. Besides, regulatory compliance for AI agents is increasingly defined by regulations such as the EU AI Act and GDPR. For instance, a comprehensive analysis of the EU AI Act's application to AI agents is provided in [139], which outlines four key governance pillars: risk assessment, transparency tools, technical deployment controls, and human oversight. It further specifies how compliance obligations, such as risk identification, agent identification, activity logging, and emergency shutdowns, should be allocated across the agent value chain. Complementing this, Chan et al. [137] highlight the importance of technical measures for data traceability, privacy protection, and auditability to support regulatory investigations and uphold user rights under GDPR.
@@
Global IoA deployments face jurisdictional fragmentation and conflicting regulations, necessitating harmonized cross-border enforcement protocols and data-sovereignty agreements. Moreover, apportioning liability in multi-agent collaborations, where agents autonomously negotiate and delegate tasks, requires novel legal constructs to assign fault among developers, operators, and the agents themselves, without stifling innovation.
3) Ethical and Responsible AI Agents. Embedding ethical reasoning and value alignment is essential for responsible AI autonomy. Pujari et al. [147] advocate for integrating ethical principles, such as virtue ethics, deontology, and consequentialism, into agent decision-making, and highlight the need for norm-aware, transparent behavior and adaptive governance. Hu et al. [148] further examine the challenges of self-sovereign decentralized agents (DeAgents) on trustless infrastructures, emphasizing the importance of participatory governance, flexible protocols, and robust fail-safes to ensure alignment with societal values even in decentralized and permissionless environments.
@@ IV Future Research Directions
IV-A IoA Standardization and Interoperability Frameworks
The widespread adoption of IoA systems necessitates unified standardization frameworks to ensure cross-platform and cross-domain interoperability. Current standardization efforts are fragmented across multiple organizations. For instance, IEEE SA-P3394111https://standards.ieee.org/ieee/3394/11377/ focuses on LLM agent interface specifications, while ITU-T F.748.46222https://www.itu.int/rec/T-REC-F.748.46-202503-I/en addresses AI agent requirements and evaluation methods. A critical challenge is the absence of unified protocols for agent discovery, capability negotiation, and dynamic service composition across heterogeneous platforms. Current approaches rely on proprietary interfaces that limit cross-platform collaboration and create vendor lock-in scenarios. Additionally, inconsistent security models and authentication mechanisms hinder trust establishment between agents from different organizations or domains. Future research should prioritize developing standardized agent description languages and lightweight communication protocols that enable seamless integration while preserving autonomy.
IV-B Secure & Adaptive Agent Communication Protocols
@@
The design of inter‐agent protocols navigates a versatility-efficiency-portability trilemma [11]: highly versatile protocols handle diverse tasks but incur heavy overhead, while efficient and portable APIs lack flexibility. Agents offer a compromise, using natural language and dynamic tool calls, but at the expense of latency, energy cost, and hallucination risks. Furthermore, as agent teams form and reconfigure dynamically, communication protocols should adapt in real time to changing roles, contexts, and capabilities. Moreover, existing MCP-based ecosystems remain susceptible to tool‐poisoning attacks whereby malicious agents inject compromised plugins or APIs; decentralized A2A networks can suffer Sybil attacks from adversary-controlled identities; and Agora's reliance on natural-language routines introduces hallucination vulnerabilities when LLM-generated schemas diverge from intended semantics. Future IoA communication should embed rigorous security guarantees to build secure agent ecosystems.
IV-C Decentralized and Self-Governing Agent Ecosystems
@@
IV-D Agent-Based Economic Systems
@@
IV-E Privacy-Preserving Agent Interactions
@@
IV-F Cyber-Physical Secure IoA